Two spherical conductors B and C having equal radii and carrying equal charges repel each other with a force F when kept at some distance. A third spherical conductor having same radius as that of B but uncharged is brought in contact with B then with C and removed away from both. What is the new force of repulsion between B and C?

Grade:12

1 Answers

Vikas TU
7 years ago
Let the initial charge on B and C is q,

After third sphere comes in contact with B then new charge on B and the third conductor would be  =q/2

and third conductor come in contact with C so total charge =q/2 + q =3q/2

From Equal distribution of charge,

Charge on C = 3q/4

 

   now force = k (q/2)(3q/4) /d2

                          = 3F/8